LASER INDUCED FLUORESCENCE SPECTRUM OF IRIDIUM MONOPHOSPHIDE

H.F. PANG, ANWEN LIU AND \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong..

Laser induced fluorescence spectrum of IrP in the spectral region between 380-600 nm has been studied. Reacting laser ablated iridium atoms with 1% PH_{3} seeded in argon produced the IrP molecule. A few vibronic transitions have been recorded. Preliminary analysis of the rotational structure indicated that these vibronic bands are with Ω′ = 0 and Ω′′ = 0 and is likely to be ^{1}Σ - X ^{1}Σ transition. Vibrational separation of the excited state is estimated to be about 442 cm^{−1}. The ground state bond length is determined to be 1.766 Å. This work represents the first experimental investigation of the spectra of IrP.